# A Pioneering Study and An Innovative Information Theory-based Approach to Enhance The Transparency in Phishing Detection

### v1

@@ Abstract.
Phishing attacks have become a serious and challenging issue for detection, explanation, and defense. Despite more than a decade of research on phishing, encompassing both technical and non-technical remedies, phishing continues to be a serious problem. Nowadays, AI-based phishing detection stands out as one of the most effective solutions for defending against phishing attacks by providing vulnerability (i.e., phishing or benign) predictions for the data. However, it lacks explainability in terms of providing comprehensive interpretations for the predictions, such as identifying the specific information that causes the data to be classified as phishing. To this end, we propose an innovative deep learning-based approach for email (the most common phishing way) phishing attack localization. Our method can not only predict the vulnerability of the email data but also automatically figure out and highlight the most important and phishing-relevant information (i.e., sentences) in each phishing email. The selected information indicates useful explanations for the vulnerability of the phishing email data. The rigorous experiments on seven real-world email datasets show the effectiveness and advancement of our proposed method in providing comprehensive explanations (by successfully figuring out the most important and phishing-relevant information in phishing emails) for the vulnerability of corresponding phishing data with higher performances from nearly (1% to 3%) and (1% to 4%) in two main Label-Accuracy and Cognitive-True-Positive measures, respectively, compared to the state-of-the-art potential baselines.
@@ 1. Introduction
The widespread adoption of artificial intelligence (i.e., using machine learning-based and deep learning-based approaches) has brought substantial influences and great success in various domain applications such as autonomous driving (Chen et al., 2023b)), data generations (Devlin et al., 2018; Radford et al., 2019; Raffel et al., 2019)), drug discovery (Paul et al., 2021)), and malware vulnerability detection (Li et al., 2018; Nguyen et al., 2021b)). By leveraging the power of machine learning and deep learning, there have been many efforts proposed for solving phishing attack problems from phishing attack detection (Ramanathan and Wechsler, 2013; Le et al., 2018; Tyagi et al., 2018; Sahingoz et al., 2019; Yang et al., 2019; Das et al., 2019; Xiao et al., 2020; Yang et al., 2021) to the investigation of the cognitive bias’s impact used in email phishing (Parsons et al., 2015; Van Der Heijden and Allodi, 2019). It has been proven that using machine learning and deep learning-based algorithms to detect phishing attacks in the early stages is one of the most effective solutions for preventing and reducing the negative effects caused. Recently, with ongoing capabilities and improvements in Natural language processing (NLP) and the popularity of large language models, e.g., GPT2 (Radford et al., 2019), BERT (Devlin et al., 2018), and T5 (Raffel et al., 2019), some of the early attempts use large language models for phishing attack-related issues have been introduced, e.g., scam-baiting mail servers that can conduct scam-baiting activities automatically (Chen et al., 2023a).
@@
”In addition to predicting the vulnerability (i.e., phishing or benign) of the data, how to derive an effective deep learning-based method that can also automatically figure out and highlight the important and phishing information in each phishing data for providing a useful and comprehensive explanation about the vulnerability of the data to users? (we name this problem as phishing attack localization)”
To this end, in the scope of our paper, we study the phishing attack localization problem on phishing emails (i.e., the most common phishing way) where we propose an innovative information theory-based model to solve the problem. Our method not only can detect the vulnerability (i.e., phishing or benign) of the email but also automatically learns and identifies the most important and phishing-relevant information (e.g., sentences) in each phishing email. The selected phishing-relevant information (e.g., sentences) helps provide useful explanations about the vulnerability of the phishing email data. It is worth noting that the ability to figure out the important and phishing-relevant information that causes the associated email data to be classified as phishing to provide a corresponding comprehensive interpretation is the main difference between phishing attack localization and phishing attack detection.
@@
  • Besides proposing a deep learning-based approach for phishing attack localization, based on the explainable machine learning and email phishing domain knowledge, we propose some appropriate measures including Label-Accuracy, Cognitive-True-Positive, and SAC (refer to Section 4.1.3 for details) for the problem.
  • We comprehensively evaluate our method on real-world phishing email datasets. The rigorous and extensive experiments show the superiority of our method over the state-of-the-art baselines.
@@ 2.1. Phishing attack detection
Machine learning has emerged as a natural choice for addressing classification challenges, including the detection of phishing attempts. Over the past decades, extensive research has been dedicated to this area. Numerous machine learning methods have been employed to tackle the issue, creating models capable of classifying content as either benign or phishing (Tyagi et al., 2018; Rao and Pais, 2019; Li et al., 2019; Zamir et al., 2020). These models leverage classification principles and carefully choose features that represent the content’s distinctive characteristics. However, machine learning-based methods are heuristics and require the knowledge of domain experts to manually engineer features from the data that can be outdated and biased. To overcome these limitations, deep learning-based solutions have been proposed for the phishing attack detection problem (Xiao et al., 2020; Abdelnabi et al., 2020; Yang et al., 2019; Lin et al., 2021) and have shown significant advantages over machine learning approaches that use hand-crafted features. AI-based phishing detection methods can predict the vulnerability label (i.e., phishing or benign) of the data (e.g., emails). However, they lack the ability to provide comprehensive explanations for their predictions.
@@ 2.2. Phishing attack localization
In our paper, we demonstrate that interpretable machine learning methods are strongly suitable for phishing attack localization in automatically learning and figuring out the most important features of the data obtained from the models to explain the model’s prediction decision. We compared the performance of our proposed method with some popular and state-of-the-art interpretable machine learning approaches including L2X ((Chen et al., 2018b)), INVASE ((Yoon et al., 2019a), ICVH ((Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)) (please refer to Section 4.1.4 in the experiments section for our brief description of these methods).
@@ 3.1. The problem statement
It is worth noting that for almost all publicly available phishing-relevant data (e.g., emails), there are only labels related to the data’s vulnerability (phishing or benign) by domain experts with the assistance of machine learning or deep learning tools. We almost do not have the ground truth of phishing information (the information truly causes the data phishing). In the phishing attack localization problem, we name this context as a weakly supervised setting where during the training process, we only use the vulnerability label of the data while not using the ground truth of phishing information existing in the data when solving the phishing attack localization problem.
@@ 3.2. Methodology
We name our proposed method as AI2TALE. In the following, we present details of how our AI2TALE  method works and addresses the phishing attack localization problem to tackle and improve the explainability (transparency) of phishing detection. An overall visualization of our method is depicted in Figure 1.
Figure 1. A visualization of our proposed AI2TALE  method for the phishing attack localization problem.
3.2.1. Learning to select important and phishing-relevant information and the training principle
@@
To make this computational process (i.e., the process consists of sampling operations from a Multi-Bernoulli distribution) continuous and differentiable during training, we apply the temperature-dependent Gumbel-Softmax trick (Eric Jang, 2016; Maddison et al., 2016) for relaxing each Bernoulli variable $z_{i}$. We sample $z_{i}\left(X;\alpha\right)\sim\text{Concrete}(\omega_{i}(X;\alpha),1-\omega_{i}(X;\alpha))$:
@@
Leveraging this property of mutual information and inspired by (Chen et al., 2018c; Nguyen et al., 2021a), we maximize the mutual information between $\tilde{X}$ and $Y$ as mentioned in Eq. (1) with the intuition is that by using the information from $Y$, the selection process $\zeta$ will be learned and enforced to obtain the most meaningful $\tilde{X}$ (i.e., $\tilde{X}$ can predict the vulnerability $Y$ of $X$ correctly).
@@
It is important to highlight that mutual information plays a crucial role in fostering collaboration between the classifier and the selection process during training. Specifically, the classifier focuses on identifying a subset of features crucial for assigning labels to data samples while the selection process is tailored to choose the optimal subset based on the feedback provided by the classifier.
@@
[…]
@@ 3.2.3. A summary of our AI2TALE  method
Algorithm 1 shows the details of our proposed AI2TALE  method in the training phase. It is worth noting that during the training process, our model is trained to learn and figure out the most important and phishing-relevant sentences in each corresponding email without using any information about the ground truth of phishing-relevant sentences. This shows a great advantage of using our method for phishing attack localization in practice because in reality, to almost all publicly available phishing-relevant email datasets, there is only information about the vulnerability $Y$ (i.e., phishing or benign) of the data by domain experts with the help of machine learning and deep learning tools.
@@
After the training phase, the selection network $\zeta$ is capable of selecting the most important and phishing-relevant sentences of a given email data $X$ by offering a high value for the corresponding coordinates $\omega_{i}\left(X;\alpha\right)$, meaning that $\omega_{i}\left(X;\alpha\right)$ represents the influence level of the sentence $\boldsymbol{x}_{i}$. We hence can pick out the most relevant sentences based on the magnitude of $\omega_{i}\left(X;\alpha\right)$. Using the selected information, the trained classifier then can predict the vulnerability of the associated email data. In our paper, with the aim of providing a highly qualified and concise explanation of the vulnerability of email data to users, we primarily assess the model’s performance based on the most important (top-1) selected sentence from each email.
@@
2       Choose a mini-batch of embedded phishing-relevant emails denoted by $\{(X_{i},Y_{i})\}_{i=1}^{m}$. Update the classifier’s parameter $\beta$ via minimizing the following cross-entropy loss $\mathcal{L}_{ce}\mathbb{E}_{X,Y}\mathbb{E}_{\mathbf{r}\sim B(0.5)}[\mathcal{L}_{ce}(Y,f_{\beta}(X\odot\mathbf{r})]$ using the Adam optimizer (Kingma and Ba, 2014). Note that minimizing this function is equivalent to maximizing the objective function in Eq. (10). Update the classifier’s parameter $\beta$ and the selection model parameter’s $\alpha$ via minimizing the following objective function $\mathbb{E}_{X,Y}\mathbb{E}_{\mathbf{z}\sim q_{sel}(\mathbf{z}|X;\alpha)}[\mathcal{L}_{ce}(Y,f_{\theta}(X\odot\mathbf{z}))]+\lambda\mathbb{E}_{X}[\sum_{i=1}^{L}\frac{\omega_{i}\left(X;\alpha\right)}{2\sigma^{2}}\left\|\boldsymbol{x}_{i}^{2}\right\|]$ using the Adam optimizer.
@@ 4.1. Experimental Designs
The fundamental goal of this experiment section is to evaluate our proposed AI2TALE  method and compare it with several potential baseline approaches for phishing attack localization. Below, we present the main research questions of our paper.
(RQ1) Can our AI2TALE  method successfully solve the phishing attack localization in terms of predicting and giving explanations in a weakly supervised setting? Can the selected information (i.e., sentences) also successfully reflect the human cognitive principles used in phishing emails?
@@
Via this research question, we mainly aim to investigate the effectiveness of our AI2TALE  method and baselines in solving the phishing attack localization problem. In particular, we want to determine their success and capability in predicting the vulnerability of email data as well as in learning and identifying the key important and phishing-relevant information (e.g., sentences) that also can reflect the human cognitive principles/triggers (i.e., the definition of ”principles of influence”, namely Reciprocity, Consistency, Social Proof, Authority, Liking, and Scarcity) used in phishing emails (e.g., (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019)).
(RQ2) Does the top-$1$ important selected sentences from our AI2TALE  method focus more on the commonly used and important cognitive principles (i.e., Scarcity, Consistency, and Authority) in phishing emails compared to those of the baselines?
As studied and pointed out in phishing attack-relevant research (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019), among cognitive triggers used in phishing attacks, some of them (Authority, Scarcity, and Consistency) are more popular than others (Reciprocity, Social, Proof, and Liking). Therefore, in this research question, we investigate how our AI2TALE  method compared to the baselines in focusing and detecting more important and popular cognitive triggers used in phishing emails through the top-$1$ important selected sentences.
(RQ3) Can our proposed information bottleneck theory training term and data-distribution mechanism be successful in improving the data representation learning and supporting the process of learning and figuring out the true phishing-relevant sentences in phishing emails?
@@
To this end, we propose two innovative solutions including the information bottleneck theory training term and data-distribution mechanism. In short, our innovative information bottle-neck training mechanism ensures that only important and label-relevant information (i.e., sentences) will be kept and selected while the data-distribution mechanism ensures the learnable selection process respecting the data distribution to select the meaningful and label-relevant information. It also aids in improving the generalization capability of the selection network, thereby enhancing the performance of the classifier through a random selection process that incorporates data augmentation (please refer to Section 3.2.2 for our detailed explanation).
In this research question, we aim to assess the effectiveness of our proposed information bottleneck theory training term (mentioned in Eq.(9)) and data-distribution mechanism (mentioned in Eq. (10)) in improving the capability of our AI2TALE  method in the process of learning and figuring out the important and true phishing-relevant information (i.e., sentences) in phishing-relevant emails.

### v2

# An Innovative Information Theory-based Approach to Tackle and Enhance The Transparency in Phishing Detection

@@ Abstract.
Phishing attacks have become a serious and challenging issue for detection, explanation, and defense. Despite more than a decade of research on phishing, encompassing both technical and non-technical remedies, phishing continues to be a serious problem. Nowadays, AI-based phishing detection stands out as one of the most effective solutions for defending against phishing attacks by providing vulnerability (i.e., phishing or benign) predictions for the data. However, it lacks explainability in terms of providing comprehensive interpretations for the predictions, such as identifying the specific information that causes the data to be classified as phishing. To this end, we propose an innovative deep learning-based approach for email (the most common phishing way) phishing attack localization. Our method can not only predict the vulnerability of the email data but also automatically learn and figure out the most important and phishing-relevant information (i.e., sentences) in the phishing email data where the selected information indicates useful and concise explanations for the vulnerability. The rigorous experiments on seven real-world diverse email datasets show the effectiveness and advancement of our proposed method in selecting crucial information, offering concise explanations (by successfully figuring out the most important and phishing-relevant information) for the vulnerability of the phishing email data. Particularly, our method achieves a significantly higher performance, ranging from approximately 1.5% to 3.5%, compared to state-of-the-art baselines, as measured by the combined average performance of two main metrics Label-Accuracy and Cognitive-True-Positive.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; September 30 - October 2, 2024; Padua, Italycopyright: acmlicensedconference: 27th International Symposium on Research in Attacks, Intrusions, and Defenses; September 30 - October 2, 2024; Padua, Italybooktitle: RAID, September 30 - October 2, 2024, Padua, Italyprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The widespread adoption of artificial intelligence (i.e., using machine learning-based and deep learning-based approaches) has brought substantial influences and great success in various domain applications such as autonomous driving (Chen et al., 2023b), data generations (Devlin et al., 2018; Radford et al., 2019; Raffel et al., 2019), drug discovery (Paul et al., 2021), and malware vulnerability detection (Li et al., 2018; Nguyen et al., 2019, 2020, 2021b, 2022b, 2022a; Fu et al., 2023)). By leveraging the power of machine learning and deep learning, there have been many efforts proposed for solving phishing attack problems from phishing attack detection (Ramanathan and Wechsler, 2013; Le et al., 2018; Tyagi et al., 2018; Sahingoz et al., 2019; Yang et al., 2019; Das et al., 2019; Xiao et al., 2020; Yang et al., 2021) to the investigation of the cognitive bias’s impact used in email phishing (Parsons et al., 2015; Van Der Heijden and Allodi, 2019). It has been proven that using machine learning and deep learning-based algorithms to detect phishing attacks in the early stages is one of the most effective solutions for preventing and reducing the negative effects caused. Recently, with ongoing capabilities and improvements in Natural language processing (NLP) and the popularity of large language models, e.g., GPT2 (Radford et al., 2019), BERT (Devlin et al., 2018), and T5 (Raffel et al., 2019), some of the early attempts use large language models for phishing attack-related issues have been introduced, e.g., scam-baiting mail servers that can conduct scam-baiting activities automatically (Chen et al., 2023a).
@@
”In addition to predicting the vulnerability (i.e., phishing or benign) of the data, how to derive an effective deep learning-based method that can also automatically learn and figure out the most important and phishing-relevant information (i.e., sentences) for providing a useful and concise explanation about the vulnerability of the phishing data to users? (in our paper, we name this problem as phishing attack localization)”
To this end, in the scope of our paper, we study the phishing attack localization problem on phishing emails (i.e., the most common phishing way) where we propose an innovative information theory-based model to solve the problem. Our method not only can detect the vulnerability (i.e., phishing or benign) of the email data but also can automatically learn and identify the most important and phishing-relevant information (e.g., sentences) in phishing emails. The selected information helps provide useful and concise explanations about the vulnerability of the phishing email data. It is worth noting that the ability to figure out the important and phishing-relevant information that causes the associated email data to be classified as phishing to provide a corresponding comprehensive interpretation is the main difference between phishing attack localization and phishing attack detection.
@@
  • Besides proposing a deep learning-based approach for phishing attack localization, based on the explainable machine learning and email phishing domain knowledge, we propose to use some appropriate measures including Label-Accuracy, Cognitive-True-Positive, and SAC (please refer to Section 4.1.3 for details) for the problem.
  • We comprehensively evaluate our method on seven real-world diverse phishing email datasets. The rigorous and extensive experiments show the superiority of our method over the state-of-the-art baselines.
@@ 2.1. Phishing attack detection
Machine learning has emerged as a natural choice for addressing classification challenges, including the detection of phishing attempts. Over the past decades, extensive research has been dedicated to this area. Numerous machine learning methods have been employed to tackle the issue, creating models capable of classifying content as either benign or phishing (Tyagi et al., 2018; Rao and Pais, 2019; Li et al., 2019; Zamir et al., 2020). These models leverage classification principles and carefully choose features that represent the content’s distinctive characteristics. However, machine learning-based methods are heuristics and require the knowledge of domain experts to manually engineer features from the data that can be outdated and biased. To overcome these limitations, deep learning-based solutions have been proposed for the phishing attack detection problem (Xiao et al., 2020; Abdelnabi et al., 2020a; Yang et al., 2019; Lin et al., 2021) and have shown significant advantages over machine learning approaches that use hand-crafted features. AI-based phishing detection methods can predict the vulnerability label (i.e., phishing or benign) of the data (e.g., emails). However, they lack the ability to provide comprehensive explanations for their predictions.
@@ 2.2. Phishing attack localization
In practice, explaining models can be divided into two categories including ”post-hoc explainability techniques” and ”intrinsic explainability techniques”. Post-hoc explainability techniques (e.g., LIME (Ribeiro et al., 2016) and SHAP (Lundberg and Lee, 2017)) aim to elucidate the decisions of a black-box model (e.g., a deep learning model, where the internal workings are not easily understandable or interpretable) without modifying the model itself. The techniques are often applied externally to the black-box model to generate explanations specific to its predictions but do not offer a comprehensive understanding of the black-box model’s internal architecture or workings. In contrast, intrinsic explainability techniques, a.k.a. self-explanatory models, (e.g., deep neural network-based methods with interpretable components such as L2X (Chen et al., 2018a) and AIM (Vo et al., 2023)) are integrated directly into a model architecture, providing interpretability as inherent features, and offering explainability as part of their design.
It is evident that intrinsic interpretable machine learning methods, a.k.a. self-explanatory models, are strongly suitable for phishing attack localization because they are not only able to make predictions themselves but also automatically learn and figure out the most important information of the data obtained from the models to explain the model’s prediction decision. In our paper, we compared the performance of our proposed method with several recent, popular, and state-of-the-art interpretable machine learning approaches falling into the category of intrinsic interpretable models including L2X (Chen et al., 2018a), INVASE (Yoon et al., 2019a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023) (please refer to Section 4.1.4 in the experiments section for our brief description of these methods).
@@ 3.1. The problem statement
It is worth noting that for almost all publicly available phishing-relevant data (e.g., emails), there are only labels related to the data’s vulnerability (phishing or benign) by domain experts with the assistance of machine learning or deep learning tools. We almost do not have the ground truth of phishing information (i.e., the information truly causes the data to be classified as phishing). In the phishing attack localization problem, we name this context as a weakly supervised setting where during the training process, we only use the vulnerability label $Y$ of the data while not requiring the ground truth of phishing information existing in the data when solving the phishing attack localization problem.
@@ 3.2. Methodology
We name our proposed method as AI2TALE . Here we present details of how our AI2TALE  method works and addresses the phishing attack localization problem to tackle and improve the explainability (transparency) of phishing detection. An overall visualization of our method is depicted in Figure 1.
Figure 1. A visualization of our proposed AI2TALE  method for solving the phishing attack localization problem.
3.2.1. Learning to select the important and phishing-relevant information and the training principle
@@
To make this computational process (i.e., the process consists of sampling operations from a Multi-Bernoulli distribution) continuous and differentiable during training, we apply the temperature-dependent Gumbel-Softmax trick (Jang et al., 2016; Maddison et al., 2016) for relaxing each Bernoulli variable $z_{i}$. We sample $z_{i}\left(X;\alpha\right)\sim\text{Concrete}(\omega_{i}(X;\alpha),1-\omega_{i}(X;\alpha))$:
@@
Leveraging this property of mutual information and inspired by (Chen et al., 2018b; Nguyen et al., 2021a), we maximize the mutual information between $\tilde{X}$ and $Y$ as mentioned in Eq. (1) with the intuition is that by using the information from $Y$, the selection process $\zeta$ will be learned and enforced to obtain the most meaningful $\tilde{X}$ (i.e., $\tilde{X}$ can predict the vulnerability $Y$ of $X$ correctly).
@@
The mutual information facilitates a joint training process for the classifier and the selection process. The classifier learns to identify a subset of features leading to a data sample’s label while the selection process is designed to select the best subset of features according to the feedback of the classifier.
@@
To make this computational procedure (i.e., it consists of sampling operations from a Multi-Bernoulli distribution) continuous and differentiable during the training process, we apply to use the temperature-dependent Gumbel-Softmax trick (Jang et al., 2016; Maddison et al., 2016) for relaxing each Bernoulli variable $r_{i}\in\mathbf{r}$ using the RelaxedBernoulli distribution function from (TensorFlowAPI, 2023). Noting that in Eq.(10), by setting $\mathbf{r}\sim\text{B}(0.5)$ to randomly select sentences from each email instead of using the entire email content, we also aim to introduce randomness into the training corpus, akin to data augmentation, which helps enhance the generalization capability of the selection network that facilitates the performance of the classifier.
@@ 3.2.3. A summary of our AI2TALE  method
Algorithm 1 shows the details of our proposed AI2TALE  method in the training phase. It is worth noting that during the training process, our model is trained to learn and figure out the most important and phishing-relevant sentences in corresponding emails without using any information about the ground truth of phishing-relevant sentences. This shows a great advantage of using our method for phishing attack localization in real-world scenarios because in practice, to almost all publicly available email datasets, there is only information about the vulnerability $Y$ (i.e., phishing or benign) of the data by domain experts with the help of machine learning and deep learning tools.
@@
After the training phase, the selection network $\zeta$ is capable of selecting the most important and phishing-relevant sentences of a given email data $X$ by offering a high value for the corresponding coordinates $\omega_{i}\left(X;\alpha\right)$, meaning that $\omega_{i}\left(X;\alpha\right)$ represents the influence level of the sentence $\boldsymbol{x}_{i}$. We hence can pick out the most relevant sentences based on the magnitude of $\omega_{i}\left(X;\alpha\right)$. Using the selected information, the trained classifier then can predict the vulnerability of the associated email data. In our paper, with the aim of providing the most highly qualified and concise explanation of the vulnerability of email data to users, we primarily assess the model’s performance based on the most important (top-1) selected sentence from each email.
@@
2       Choose a mini-batch of embedded phishing-relevant emails denoted by $\{(X_{i},Y_{i})\}_{i=1}^{m}$. Update the classifier’s parameter $\beta$ via minimizing the following cross-entropy loss $\mathcal{L}_{ce}\mathbb{E}_{X,Y}\mathbb{E}_{\mathbf{r}\sim B(0.5)}[\mathcal{L}_{ce}(Y,f_{\beta}(X\odot\mathbf{r})]$ using the Adam optimizer (Kingma and Ba, 2014). Note that minimizing this function is equivalent to maximizing the objective function in Eq. (10).
3       Update the classifier’s parameter $\beta$ and the selection model parameter’s $\alpha$ via minimizing the following objective function $\mathbb{E}_{X,Y}\mathbb{E}_{\mathbf{z}\sim q_{sel}(\mathbf{z}|X;\alpha)}[\mathcal{L}_{ce}(Y,f_{\theta}(X\odot\mathbf{z}))]+\lambda\mathbb{E}_{X}[\sum_{i=1}^{L}\frac{\omega_{i}\left(X;\alpha\right)}{2\sigma^{2}}\left\|\boldsymbol{x}_{i}^{2}\right\|]$ using the Adam optimizer.
4
@@ 4.1. Experimental Designs
The fundamental goal of this experiment section is to evaluate our proposed AI2TALE  method and compare it with several state-of-the-art baseline approaches for phishing attack localization. Below, we present the main research questions of our paper.
(RQ1) Can our AI2TALE  method solve the phishing attack localization by predicting and giving explanations in a weakly supervised setting? Does the selected information (i.e., sentences) also reflect the human cognitive principles used in phishing emails?
@@
Via this research question, we mainly aim to investigate the effectiveness of our AI2TALE  method and baselines in solving the phishing attack localization problem. In particular, we evaluate their capability in predicting each email’s vulnerability and identifying the most important (top-1) phishing-relevant sentence for offering the most concise and highly qualified explanation. The selected sentence in each email can also reflect the human cognitive principles/triggers (i.e., the definition of ”principles of influence”, namely Reciprocity, Consistency, Social Proof, Authority, Liking, and Scarcity) often used in phishing emails (e.g., (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019)).
(RQ2) Does the most important (top-$1$) selected sentences from our AI2TALE  method primarily consist of the commonly used and important cognitive principles (i.e., Scarcity, Consistency, and Authority) in phishing emails compared to those of the baselines?
As studied and pointed out in phishing attack-relevant research (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019), among cognitive triggers used in phishing attacks, some of them (Authority, Scarcity, and Consistency) are more popular than others (Reciprocity, Social, Proof, and Liking). Therefore, in this research question, we investigate how our AI2TALE  method compared to the baselines in consisting of more important and popular cognitive triggers used in phishing emails in the most important (top-$1$) selected sentences.
(RQ3) Can our proposed information bottleneck theory training term and data-distribution mechanism be successful in improving the data representation learning and supporting the process of learning and figuring out the most important and phishing-relevant sentences in phishing emails?
@@
To this end, we further propose two innovative solutions including the information bottleneck theory training term and data-distribution mechanism. In short, our innovative information bottle-neck training mechanism ensures that only important and label-relevant information (i.e., sentences) will be kept and selected while the data-distribution mechanism ensures the learnable selection process respecting the data distribution to select the meaningful and label-relevant information. It also aids in improving the generalization capability of the selection network, thereby enhancing the performance of the classifier through a random selection process that incorporates data augmentation (please refer to Section 3.2.2 for our detailed explanation).
In this research question, we aim to assess the effectiveness of our proposed information bottleneck theory training term (mentioned in Eq.(9)) and data-distribution mechanism (mentioned in Eq. (10)) in improving the capability of our AI2TALE  method in the process of learning and figuring out the most important and truly phishing-relevant information (i.e., sentences) in phishing-relevant emails.